### v2

@@ 4.6 Local convergence of function values
Taking $p>1$ in the proof of Corollary 4.30, linear convergence rates could be obtained as in Corollary 4.26 for the iterates.
4.7 Weak convergence
Example 4.33.
Suppose $M=A^{*}A$ for some $A\in\mathbb{L}(X;V)$ for a Hilbert space $V$. Then the $M$-semi-norm convergence clearly implies $Ax^{k}\to 0$, and consequently $Mx^{k}\to 0$. Thus $M$ is weak-* admissibility. In Hilbert spaces, every positive semi-definite self-adjoint operator has such a square root $A$ with $V=X$. For a convolution-based construction in the space of Radon measures, see [34, Theorem 2.4].
Theorem 4.34.
Suppose Sections 4.2 and 4.2 hold with $p=1$ and $\eta>0$ at some $\bar{x}=\hat{x}\in H^{-1}(0)$, and that either Section 4.2Item (a) or Item (b) (only the item, not the entire assumption) holds with $\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$ and $\sum_{k=0}^{\infty}\varepsilon_{k}(\hat{x})<\infty$ at all $\hat{x}\in\hat{X}:=H^{-1}(0)\cap\mathbb{O}_{M}(\bar{x},\delta)$. Also suppose that the preconditioner $M$ is admissible for weak convergence, and $F$ is either convex or $F^{\prime}$ is weak-to-strong continuous. Then $x^{k}\mathrel{\rightharpoonup}\hat{x}$ weakly for some $\hat{x}\in\hat{X}$.
Proof 4.35.
Lemma 4.24 proves that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, as well as that $\sup_{N\in N}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$. The latter establishes $\|x^{k+1}-x^{k}\|_{M}\to 0$, and through admissibility for weak convergence, and (28), that ~$\tilde{\partial}_{k+1}=-M(x^{k+1}-x^{k})\to 0$ strongly in $X^{*}$. Moreover, Section 4.2 yields ~$\|\tilde{\partial}_{k+1}-x^{*}_{k+1}\|_{X^{*}}\to 0$ for some $x^{*}_{k+1}\in H(x^{k+1})$. Consequently $x^{*}_{k+1}\to 0$. Since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$, as in the proof of Lemma 4.24, we show the quasi-Féjer monotonicity (48) for all $\hat{x}\in\hat{X}$ and $k\in\mathbb{N}$.
  1. 1.
    If $F$ is convex, $H$ is maximally monotone [^4] That the additive skew-adjoint term $\Xi$ does not destroy maximal monotonicity, can be proved completely analogously to the Hilbert space case in [7, Lemma 9.9]., hence weak-to-strong outer semicontinuous [7, Lemma 6.10]. Now $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$ and $H(x^{k_{j}+1})\ni x^{*}_{k_{j}+1}\to 0$ obliges $0\in H(\hat{x})$.
  2. 2.
    Suppose then that $F^{\prime}$ is weak-to-strong continuous. Now still $P:x\mapsto\partial G(x)+\Xi x$ is maximally monotone4, hence weak-to-strong outer semicontinuous. We have $P(x^{k_{j}+1})\ni x^{*}_{k_{j}+1}-F^{\prime}(x^{k_{j}+1})\to-F^{\prime}(\hat{x})$ strongly in $X^{*}$, as well as $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$, so we must have $-F^{\prime}(\hat{x})\in P(\hat{x})$. But this again says $0\in H(\hat{x})$.
Thus every weak limiting point $\hat{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$ satisfies $0\in H(\hat{x})$. But, since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, also $\hat{x}\in\mathbb{O}_{M}(\bar{x},\delta)$. This proves that $\hat{x}\in\hat{X}$. Since, by assumption, $\sum_{k=0}^{\infty}\varepsilon_{k}(\hat{x})<\infty$ for all $\hat{x}\in\hat{X}$, the quasi-Féjer monotonicity (48) with the quasi-Opial’s Lemma A.2 finishes the proof.
Example 4.36.
In the setting of Section 2 and Theorem 2.16, the weak-*-to-strong continuity of $F^{\prime}$ can be achieved, for example, when $F(x)=\frac{1}{2}\|S(x)-b\|^{2}$ for a Lipschitz and bounded $S$ with finite-dimensional range.
Remark 4.37.
All of our theory also applies when $X$ is the dual space of a separable normed space $X_{*}$, and we replace in our definitions $X^{*}$ by the predual space $X_{*}$, that is, subdifferentials are subsets of $X_{*}$, and $M,\Lambda\in\mathbb{L}(X;X_{*})$, etc. With this change the theory applies, for example, to $X$ a space of Radon measures, as in [34]. Then Theorem 4.34 proves the weak-* convergence.
Appendix A Opial’s lemma for quasi-Féjer monotonicity

### v3

@@ 6.3 Primal-dual proximal splitting
Taking $p>1$ in the proof of Corollary˜6.16, linear convergence rates could be obtained as in Corollary˜4.12 for the iterates.
6.4 Further topics
Theorem 6.19 (PDPS with adjoint mismatch).
Assume the setup of Example˜4.1 with $\tau\sigma\|K\|^{2}\leq 1$ and, for simplicity, $f=0$ and Hilbert $Z$ and $Y$. Suppose $\operatorname{dom}h_{*}$ is bounded, and that $g$ and $h_{*}$ are, respectively, $\gamma_{g}$- and $\gamma_{h_{*}}$-strongly convex for some $\gamma_{g}>0$ and $\gamma_{h_{*}}\geq 0$. Let $\gamma:=\min\{\gamma_{g}\tau/4,\gamma_{h_{*}}\sigma/2\}$. In the PDPS (24), not able to compute $K^{*}$, replace it with a “mismatched” adjoint $K^{*\approx}$. Then, for any $\bar{x}\in Z\times Y$ and $p\in(1,1+2\gamma]$, Section˜4.2Item˜(a) holds with $\breve{\Lambda}=0$, $\Omega_{\bar{x}}=Z\times Y$, $\delta=\infty$, $r_{p}\leq\varepsilon/(1-p)$, and
Proof 6.20.
With $M$, $G$, and $F$ given by Example˜4.1, the abstract algorithm (22) reads
$M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}=x_{k+1}^{*}+((K^{*\approx}-K^{*})y^{k},0)\quad\text{for a}\quad x_{k+1}^{*}\in H(x^{k+1}).$
Here $H$ is defined in (21). Using Lemma˜6.11 in the final step, we estimate
$\langle\tilde{\partial}_{k+1}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ $=\langle\tilde{\partial}_{k+1}-x_{k+1}^{*}|x^{k+1}-\bar{x}\rangle_{X^{*},X}+\langle x_{k+1}^{*}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$
$\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Therefore, (27) holds with the stated choices. Moreover, we have $\sum_{k=0}^{N-1}p^{k-N}\leq 1/(p-1)$ for any $p\in(1,1+2\gamma]$, verifying (29) and consequently Section˜4.2Item˜(a).
Remark 6.21 (Stochastic optimisation methods).
Also stochastic optimisation methods can be approached in our framework through lifting: we take $X$ as a space of random variables on a space $\mathscr{X}$, set $F(x)=\mathbb{E}[\mathscr{F}\circ x]$ for a function $\mathscr{F}$ on $\mathscr{X}$, $Mx:=\mathbb{E}^{*}\mathscr{M}\mathbb{E}[x]$ for $\mathscr{M}\in\mathbb{L}(\mathscr{X};\mathscr{X}^{*})$, etc., where $\mathbb{E}$ is the expectation. Without inexactness, this produces parallel copies of the very same optimisation method for every random event $\omega$. We then model the random choices made on every step through inexactness, subject to Sections˜4.2, 4.2 and 4.2.
Appendix A Scalar tracking results
We prove a a simple scalar tracking result, which will be used to establish the results of Section˜2. The following is a scalar version of Section˜2.1.
Assumption A.0.
For a given $k\geq 1$, for all $j=1,\ldots,k$, the scalars ${d^{u}_{j}},{d^{w}_{j}},{\varrho_{j}}\geq 0$ and $\widetilde{d}_{j}$$\in\mathbb{R}$ satisfy for some $\pi_{u},\pi_{w},\mu_{u}>0$, and $\kappa_{u},\kappa_{w}>1$, as well as $\alpha_{u},\alpha_{w}>0$, the inequalities
(i) $\kappa_{u}{d^{u}_{j+1}}$ $\leq{d^{u}_{j}}+\pi_{u}{\varrho_{j}},$
(ii) $\kappa_{w}{d^{w}_{j+1}}$ $\leq{d^{w}_{j}}+{\mu_{u}}{d^{u}_{k+1}}+\pi_{w}{\varrho_{j}},\quad\text{and}$
(iii) ${\widetilde{d}_{j}}$ $\leq\alpha_{u}{d^{u}_{j+1}}+\alpha_{w}{d^{w}_{j+1}}.$
We wish to develop simple bounds for $\widetilde{d}_{k}$} that can be readily used in convergence proofs when Appendix˜A is instantiated as Section˜2.1. These core estimates then allow us to isolate the contributions of initialisation and update errors, and thereby quantify the impact of inexact inner and adjoint solutions over multiple iterations on the differential approximations. We start with a result that unrolls the recursion in Appendix˜A.
Lemma A.1.
Let Appendix˜AEqs.˜i and ii hold for a $k\geq 1$. Then, letting $\iota_{k}:=\sum_{m=1}^{k}\kappa_{u}^{-m}\kappa_{w}^{-(k+1-m)}$ (understanding that $\iota_{0}=0$), for all $k\geq 0$ and $\alpha_{u},\alpha_{w}\geq 0$, we have
(62) $R^{k+1}(\alpha_{u},\alpha_{w})=\alpha_{u}{d^{u}_{k+1}}+\alpha_{w}{d^{w}_{k+1}}$ $\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-k}{d^{w}_{1}}$
$+\sum_{j=0}^{k-1}\bigl(\alpha_{u}\kappa_{u}^{-(k-j)}\pi_{u}+\alpha_{w}[\iota_{k-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(k-j)}\pi_{w}]\bigr){\varrho_{j+1}}.$
Proof A.2.
For $k=1$, we have $\iota_{1}=\kappa^{-1}_{u}\kappa^{-1}_{w}$ and, by assumption, ${d^{u}_{2}}\leq\kappa^{-1}_{u}{d^{u}_{1}}+\kappa^{-1}_{u}\pi_{u}{\varrho_{1}}\text{ and }{d^{w}_{2}}\leq\kappa^{-1}_{w}{d^{w}_{1}}+\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{2}}+\kappa^{-1}_{w}\pi_{w}{\varrho_{1}}.$ Multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, then summing up, observing to cancel the two instances of $\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{2}}$, establishes (62).
We then take $k=n+1$, and proceed by induction, assuming (62) to hold for $k=n$. Again, ${d^{u}_{n+2}}\leq\kappa^{-1}_{u}{d^{u}_{n+1}}+\kappa^{-1}_{u}\pi_{u}{\varrho_{n+1}}\text{ and }{d^{w}_{n+2}}\leq\kappa^{-1}_{w}{d^{w}_{n+1}}+\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{n+1}}+\kappa^{-1}_{w}\pi_{w}{\varrho_{n+1}}$ by assumption. As in the case $k=1$, multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, and then summing up, yields
$R^{n+2}(\alpha_{u},\alpha_{w})=\alpha_{u}{d^{u}_{n+2}}+\alpha_{w}\kappa^{-1}_{w}{d^{w}_{n+2}}$ $\leq(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}}){d^{u}_{n+1}}+\alpha_{w}\kappa^{-1}_{w}{d^{w}_{n+1}}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}]){\varrho_{n+1}}.$
$R^{n+2}(\alpha_{u},\alpha_{w})$ $\leq((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-n}+\alpha_{w}\kappa^{-1}_{w}\iota_{n}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa^{-1}_{w}\kappa_{w}^{-n}{d^{w}_{1}}$
$+\sum_{j=0}^{n-1}\bigl((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-(n-j)}\pi_{u}+\alpha_{w}\kappa^{-1}_{w}[\iota_{n-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(n-j)}\pi_{w}]\bigr){\varrho_{j+1}}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}]){\varrho_{n+1}}$
$=(\alpha_{u}\kappa_{u}^{-(n+1)}+\alpha_{w}\mu_{u}(\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n})){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-(n+1)}{d^{w}_{1}}$
$+\sum_{j=0}^{n}\bigl(\alpha_{u}\kappa_{u}^{-(n+1-j)}\pi_{u}+\alpha_{w}[(\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}){\mu_{u}}\pi_{u}+\kappa_{w}^{-(n+1-j)}\pi_{w}]\bigr){\varrho_{j+1}}.$
Thus we obtain (62) for $k=n+1$.
Thus, by sum formulae for arithmetic-geometric progressions [14, formula 0.113],
The next lemma bounds the distance between the true differential and estimate at iteration $k$ through an error term. In the two lemmas that follow, we further estimate the error term. We use the first lemma directly in Theorem˜2.8.
Lemma A.3.
Suppose Appendix˜A holds for a $k\geq 1$. Then for any $p\in(0,\kappa)$, we have
(66) ${\widetilde{d}_{k}^{2}}\leq\breve{e}_{p,k},$
(68) $\breve{e}_{p,k}$ $:=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}p^{k}}{(d^{u}_{1})^{2}}+\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}p^{k}}{(d^{w}_{1})^{2}}+\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}{\varrho_{j+1}}.$
Proof A.4.
We prove for all $s\in\mathbb{R}$ that
(69) $4s{\widetilde{d}_{k}}-4s^{2}\leq\breve{e}_{p,k}.$
Maximising the left hand side with respect to $s$ then establishes (66).
By Appendix˜AEq.˜iii, we have
${\widetilde{d}_{k}}\leq\alpha_{u}{d^{u}_{k+1}}+\alpha_{w}{d^{w}_{k+1}}=:R^{k+1}.$
Since $\{x^{n}\}_{n=0}^{k}\subset\Omega$, invoking the inner and adjoint tracking Appendix˜AEqs.˜i and ii and Lemma˜A.1, we obtain
$R^{k+1}\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-k}{d^{w}_{1}}+\sum_{j=0}^{k-1}\psi_{k-j}{\varrho_{j+1}}.$
Using Young’s inequality several times here, we deduce for any $\theta_{k}^{u},\theta_{k}^{w},\theta_{k,j}>0$ that
(70) $\begin{split}4sR^{k+1}&\leq\frac{(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})^{2}}{\theta_{k}^{u}}{(d^{u}_{1})^{2}}+\frac{(\alpha_{w}\kappa_{w}^{-k})^{2}}{\theta_{k}^{w}}{(d^{w}_{1})^{2}}\\ &+\sum_{j=0}^{k-1}\frac{\psi_{k-j}^{2}}{\theta_{k,j}}{\varrho_{j+1}^{2}}+4\left(\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}\right)s^{2}.\end{split}$
Take $\theta_{k}^{u}=p^{k}\varsigma_{p}^{-1}\pi_{u}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})$, $\theta_{k}^{w}=p^{k}\varsigma_{p}^{-1}\pi_{w}\alpha_{w}\kappa_{w}^{-k}$, and $\theta_{k,j}=\varsigma_{p}^{-1}p^{k-j}\psi_{k-j}$. Observe from (63) that $\iota_{k}\leq\kappa_{w}\iota_{k+1}$. Hence $p^{k}\iota_{k}\leq(\kappa_{w}/p)p^{k+1}\iota_{k+1}$, and further, $p^{k}\psi_{k}\leq(\overline{\kappa}/p)p^{k+1}\psi_{k+1}$, where $\overline{\kappa}/p>1$. Now
Thus rearranging (70) establishes (69). Finally, the bound in (67) on $\varsigma_{p}$ follows from Eq.˜65 and $\sum_{j=0}^{\infty}(p/\kappa)^{j}=1/(1-p/\kappa)=\kappa/(\kappa-p)$.
We use the next result in Theorem˜2.5.
Lemma A.5.
Suppose Appendix˜A holds for a $k\geq 1$. Then for any $p\in[1,\kappa)$, we have
(71) ${\widetilde{d}_{k}^{2}}\leq\varsigma_{p}^{2}{\varrho_{k+1}}+e_{p,k},$
where, for $\breve{e}_{p,k}$ defined (68),
(72) $e_{p,k}:=\breve{e}_{p,k}-\varsigma_{p}^{2}{\varrho_{k+1}}.$
satisfies
(73) $\begin{split}\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}&\leq\frac{{(d^{u}_{1})^{2}}}{\pi_{u}}\bigg(\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg)+\frac{{(d^{w}_{1})^{2}}}{\pi_{w}}\bigg(\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg).\end{split}$
Proof A.6.
We only need to prove (73); the rest is immediate from Lemma˜A.3. Let
$A_{k}$ $=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}}{(d^{u}_{1})^{2}},$ $B_{k}$ $=\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}}{(d^{w}_{1})^{2}},$
$C_{k}$ $=\sum_{j=0}^{k-1}p^{j}\varsigma_{p}\psi_{k-j}{\varrho_{j+1}},\quad\text{and}$ $D_{k}$ $=p^{j}\varsigma_{p}^{2}{\varrho_{k+1}}.$
Then $p^{k}e_{p,k}=:A_{k}+B_{k}+C_{k}-D_{k}$ from Eqs.˜72 and 68. By the assumption $p\in[1,\kappa)$, we have $\frac{\overline{\kappa}}{p}p^{j}\geq 1$. Hence from (67) we have that $\varsigma_{p}=\frac{\overline{\kappa}}{p}\sum_{j=0}^{\infty}p^{j}\psi_{j}\geq\sum_{j=0}^{\infty}\psi_{j}.$ Now
$\begin{split}\sum_{k=0}^{N-1}C_{k}&=\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}p^{j}\varsigma_{p}\psi_{k-j}{\varrho_{j+1}^{2}}=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{k=j+1}^{N-1}\psi_{k-j}{\varrho_{j+1}^{2}}\\ &=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}{\varrho_{j+1}^{2}}\leq\sum_{j=0}^{N-2}p^{j}\varsigma_{p}^{2}{\varrho_{j+1}^{2}}\leq\sum_{k=0}^{N-1}D_{k}.\end{split}$
Moreover, using (65) and the sum formula for geometric series, we estimate that $\sum_{k=0}^{N-1}(A_{k}+B_{k})$ is less than the right-hand side of (73).
We use the next lemma in Theorem˜2.8.
Lemma A.7.
Suppose Appendix˜A holds for all $k\geq 1$. Then, for $\breve{e}_{1,k}$ given in (67) we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\breve{e}_{1,k}<\infty\quad\text{whenever}\quad\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{\varrho_{j+1}^{2}}<\infty.$
Proof A.8.
Observe that
(74) $\breve{e}_{1,k}=\frac{\varsigma_{1}}{\pi_{u}}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}\mu_{u}){(d^{u}_{1})^{2}}+\frac{\varsigma_{1}}{\pi_{w}}\alpha_{w}\kappa_{w}^{-k}{(d^{w}_{1})^{2}}+\sum_{j=0}^{k-1}\varsigma_{1}\psi_{k-j}{\varrho_{j+1}^{2}}.$
Using (65) and the sum formula for geometric series, we bound the sum of the first two terms over $k\in\{0,\ldots,N-1\}$ by
${(d^{u}_{1})^{2}}\frac{\varsigma_{1}}{\pi_{u}}\bigg(\frac{\alpha_{u}\kappa}{\kappa-1}+\frac{\alpha_{w}\mu_{u}}{(\kappa-1)^{2}}\bigg)+{(d^{w}_{1})^{2}}\frac{\varsigma_{1}}{\pi_{w}}\bigg(\frac{\alpha_{w}\kappa}{\kappa-1}\bigg).$
Lemma˜A.3 establishes $\varsigma_{1}=\overline{\kappa}\sum_{j=0}^{\infty}\psi_{j}<\infty.$ With this, we sum the third term of (74) over $k\in\{0,\ldots,N-1\}$, and change the order of summation to obtain
$\varsigma_{1}\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}\psi_{k-j}{\varrho_{j+1}^{2}}=\varsigma_{1}\sum_{j=0}^{N-2}\left(\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}\right){\varrho_{j+1}^{2}}\leq\frac{\varsigma_{1}^{2}}{\overline{\kappa}}\sum_{j=0}^{N-2}{\varrho_{j+1}^{2}}.$
Thus $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\breve{e}_{1,k}<\infty$ whenever $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{\varrho_{j+1}^{2}}<\infty$.
Appendix B Opial’s lemma for quasi-Féjer monotonicity